find electric field of all regions


+sigma =2c ------------------


E=3


-sigma =2c -------------------

1
Expert's answer
2021-12-01T19:19:09-0500

The diagram and values given are not clear.

Modifying the conditions of the question:


QUESTION:

Find the electric field of all regions ( I, II and III)





SOLUTION


The electric field at each region are shown in the figure below:





REGION I


At the region, the electric fields are equal in magnitude but face opposite direction.

"E_T=-E+E=0"

Hence net electric field in region I is zero


REGION II


At region 2, the electric field is uniform and face the same direction.

"E_T= E+E= 2E\\But\>E=\frac{\sigma}{2\epsilon_0}"


Net electric field in region II

"=2E=2×\frac{\sigma}{2\epsilon_0}\\=\frac{\sigma}{\epsilon_0}"


REGION III


Region III is similar to region I and the net electric field is Zero.